ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 215, 2016

Adams, T. M. and C. E. Silva Reprint of “Weak rational ergodicity does not imply rational ergodicity”, Israel Journal of Mathematics 214 (2016), 491–505

1027

Alfonseca, M. A., M. Cordier and D. Ryabogin On bodies with directly congruent projections and sections

765

Bailey, A., M. Finn-Sell and R. Snocken Subsemigroup, ideal and congruence growth of free semigroups

459

Barbatis, G. and P. D. Lamberti Monotonicity, continuity and differentiability results for the L^{p} Hardy constant

1011

Bourgain, J. A Quantitative Oppenheim theorem for generic diagonal quadratic forms

503

Bradford, H. Expansion, random walks and sieving in SL_{2}(𝔽_{p}[t])  559

Broderick, R., V. Cyr and B. Kra Complexity and directional entropy in two dimensions

135

Cannon, J., M. Garonzi, D. Levy, A. Maróti and I. I. Simion Groups equal to a product of three conjugate subgroups

31

Caranti, A. Erratum to “A module-theoretic approach to abelian automorphism groups”, Israel Journal of Mathematics 205 (2015), 235–246

1025

Casolo, C., S. Dolfi, E. Pacifici and L. Sanus Groups whose character degree graph has diameter three

523

Chandgotia, N. and T. Meyerovitch Markov random fields, Markov cocycles and the 3-colored chessboard

909

Chen, H. and C. Xi Dominant dimensions, derived equivalences and tilting modules

349

Cohen, M. P. and R. R. Kallman Openly Haar null sets and conjugacy in Polish groups

1

Cordier, M. See M. A. Alfonseca

Csikvári, P. Matchings in vertex-transitive bipartite graphs  99

Cyr, V. See R. Broderick Doležal, N., D. Priess and M. Zelený Infinite games and σ-

porosity

441

Dolfi, S. See C. Casolo Eleftherakis, G. K., R. H. Levene and I. G. Todorov Schur

idempotents and hyperreflexivity 317 Finn-Sell, M. See A. Bailey Garonzi, M. See J. Cannon Herbin, E. and A. Richard Local Hölder regularity for set-indexed

processes

397

Ibarlucía, T. The dynamical hierarchy for Roelcke precompact Polish groups

965

Iovanov, M. C., Z. Mesyan and M. L. Reyes Infinite-dimensional diagonalization and semisimplicity

801

Juhász, I., L. Soukup and Z. Szentmiklóssy Pinning down versus density

583

Kakariadis, E. T.A. A note on the gauge invariant uniqueness theorem for C^{∗}-correspondences

513

Kallman, R. R. See M. P. Cohen Kra, B. See R. Broderick Lamberti, P. D. See G. Barbatis Levene, R. H. See G. K. Eleftherakis Levy, D. See J. Cannon

Lewis, M. L. Variations on average character degrees and p-nilpotence  749

Maróti, A. See J. Cannon Mesyan, Z. See M. C. Iovanov Meyerovitch, T. See N. Chandgotia Nielsen, P. P. The behavior of ascending chain conditions on sub-

modules of bounded finite generation in direct sums

339

Priess, D. See M. Doležal Pacifici, E. See C. Casolo Ran, Z. Structure of the cycle map for Hilbert schemes of families of

nodal curves

669

Reyes, M. L. See M. C. Iovanov Richard, A. See E. Herbin Rosso, G. Derivative at s = 1 of the p-adic L-function of the symmet-

ric square of a Hilbert modular form

255

Ryabogin, D. See M. A. Alfonseca Saghin, R. and J. Yang Continuity of topological entropy for per-

turbation of time-one maps of hyperbolic flows

857

Sanus, L. See C. Casolo Silva, C. E. See T. M. Adams Simion, I. I. See J. Cannon Skrypnik, I. I. Removability of isolated singularities for anisotropic

elliptic equations with gradient absorption 163 Snocken, R. See A. Bailey Soukup, L. See I. Juhász Szentmiklóssy, Z. See I. Juhász

Thomine, D. Local time and first return time for periodic semi-flows  53

Todorov, I. G. See G. K. Eleftherakis Topaz, A. Abelian-by-central Galois groups of fields II: definability of

inertia/decomposition groups

713

Trang, N. PFA and guessing models  607

Xi, C. See H. Chen Yang, J. See R. Saghin Zelený M. See M. Doležal

Zhao, L. Rational lines on cubic hypersurfaces  877

Zhu, J. Gluing formula of real analytic torsion forms and adiabatic limit 181